# Second- and third-order properties of multidimensional Langevin equations

### v16

@@ 4. Third-order properties
(100) 00 0 $=-\int\mathrm{d}\mathbf{x}\,x^{i}x^{j}x^{k}\frac{\partial J^{l}(\mathbf{x})}{\partial x^{l}}$
@@ 5.2. Two-dimensional example
(124) $-\lambda_{x}\langle xy^{4}\rangle=-4b^{xy}_{y}\langle y^{4}\rangle.$
@@ 5.3. Integrated variables
We note that if $D^{ky}=0$ for $k$ stationary, which is a condition for detailed balance, then $\langle x^{k}(\tau)(y(\tau)-y(0))\rangle=0$. If in addition $02b^{iy}_{k}C^{jk}+2b^{jy}_{k}C^{ik}=$ 0, then $\langle x^{i}(\tau)x^{j}(\tau)(y(\tau)-y(0))\rangle=0$.
@@ 6.2. One-dimensional case
(asterisk denoting the complex conjugate), which can be easily shown using Itô’s lemma (with suitable assumptions on $f,g$). When $a=b=0$, the Koopman eigenfunctions are the probabilist’s Hermite polynomials ${He}_{n}(x)$ with eigenvalues $-n$. In the present case, the Koopman eigenfunctions $f_{n}(x)$ and the corresponding eigenvalues $\lambda_{n}$ are:
@@ 6.3. Multidimensional case
×$\qquad{}\times\left({a^{i^{\prime}}_{lm}C^{j^{\prime}l}C^{k^{\prime}m}+a^{j^{\prime}}_{lm}C^{i^{\prime}l}C^{k^{\prime}m}+a^{k^{\prime}}_{lm}C^{i^{\prime}l}C^{j^{\prime}m}+b^{i^{\prime}j^{\prime}}_{l}C^{k^{\prime}l}+b^{i^{\prime}k^{\prime}}_{l}C^{j^{\prime}l}+b^{j^{\prime}k^{\prime}}_{l}C^{i^{\prime}l}}\right)$
@@ 6.4. Generalized Koopman eigenfunctions
$\gamma^{ij}_{klm}$ :=$:=\left\{{\left[{(\mathbf{A}\otimes\mathds{1}+\mathds{1}\otimes\mathbf{A})\otimes\mathds{1}\otimes\mathds{1}\otimes\mathds{1}-\mathds{1}\otimes\mathds{1}\otimes(\mathbf{A}\otimes\mathds{1}\otimes\mathds{1}+\mathds{1}\otimes\mathbf{A}\otimes\mathds{1}+\mathds{1}\otimes\mathds{1}\otimes\mathbf{A})}\right]^{-1}}\right\}^{ijk^{\prime}l^{\prime}m^{\prime}}_{i^{\prime}j^{\prime}klm}$
(173) ×$\quad{}\times\left({a^{i^{\prime}}_{k^{\prime}l^{\prime}}\delta^{j^{\prime}}_{m^{\prime}}+a^{j^{\prime}}_{k^{\prime}l^{\prime}}\delta^{i^{\prime}}_{m^{\prime}}}\right).$
@@
×$\qquad{}\times\left({x^{k^{\prime}}(0)x^{l^{\prime}}(0)x^{m^{\prime}}(0)-C^{k^{\prime}l^{\prime}}x^{m^{\prime}}(0)-C^{k^{\prime}m^{\prime}}x^{l^{\prime}}(0)-C^{l^{\prime}m^{\prime}}x^{k^{\prime}}(0)}\right)$
@@ 6.7. Finite time interval inhomogeneous diffusion
However, for finite $\tau$, the above quantity can be non-zero for homogeneous diffusion because of non-vanishing third moments. This can be illustrated by the example of Eq. (93) in one dimension with $-A=D=1$ and $b=0$. From Eqs. (156)–(157), we have:
@@ 6.8. Quantitative significance of third-order effects
×$\qquad\qquad\qquad{}\times\left({D^{yz}-A^{y}_{p}(\mathbf{A}^{-1})^{p}_{q}D^{qz}-A^{z}_{r}(\mathbf{A}^{-1})^{r}_{s}D^{sy}+A^{y}_{p}A^{z}_{r}(\mathbf{A}^{-1})^{p}_{q}(\mathbf{A}^{-1})^{r}_{s}D^{qs}}\right).$
@@
.$\qquad\Big{.}{{}+C^{ji^{\prime}}(D^{ij^{\prime}}D^{kk^{\prime}}-D^{ik^{\prime}}D^{kj^{\prime}})+C^{jj^{\prime}}(D^{ii^{\prime}}D^{kk^{\prime}}-D^{ik^{\prime}}D^{ki^{\prime}})}\Big{]}.$
@@ 6.10. Comparison of experimental data and theoretical predictions
We consider the sum and difference of these quantities at $\tau=0$. The former is equal to $(\Delta t)L(x^{i}x^{j},y)$, while the latter is equal to $-2(\Delta t)(C^{ik}b^{jy}_{k}+C^{jk}b^{iy}_{k})$. We can calculate the “ensemble variances” of these quantities using the aforementioned prescriptions. We may consider using them for comparison at arbitrary $\tau$. However, we may be interested in comparing the above quantities directly. They may be obtained by taking one-half the sum and difference of the sum and difference. We may treat the sum and difference of the above quantities as uncorrelated in the “ensemble covariance” as they acquire opposite signs under time reversal. Thus we conclude that the above quantities should be compared according to one-fourth the sum of the “ensemble variances” of $(\Delta t)L(x^{i}x^{j},y)$ and $-2(\Delta t)(C^{ik}b^{jy}_{k}+C^{jk}b^{iy}_{k})$. As in the linear case, and as is analogously true for all quantities, if the theoretical values at $\tau=0$ exceed the prescriptions of the “ensemble variances” (for either second- or third-order quantities), we may need to increase the tolerance for comparison accordingly.
@@ 7.2. Detailed balance
Here we discuss the conditions for detailed balance in the Fokker–Planck equation Eq. (97) in the presence of odd variables. Each variable $x^{i}$ has a parity $\epsilon^{i}=\pm 1$, where under time-reversal $x^{i}\to\epsilon^{i}x^{i}$ ($i$ not summed over), and the product $\bm{\epsilon}\mathbf{x}$ is defined as $(\bm{\epsilon}\mathbf{x})^{i}=\epsilon^{i}x^{i}$ ($i$ not summed over), following [3]. In terms of the stationary distribution $p(\mathbf{x})$, the conditions for detailed balance are [3]:
@@ 7.3. Inhomogeneous diffusion
(248) $\mathbf{x}^{(2)+}$ :=$:=\{{x^{+}}^{i}{x^{+}}^{j}-\langle{x^{+}}^{i}{x^{+}}^{j}\rangle,{x^{-}}^{i}{x^{-}}^{j}-\langle{x^{-}}^{i}{x^{-}}^{j}\rangle\}_{i,j}$
(249) $\mathbf{x}^{(2)-}$ :=$:=\{{x^{+}}^{i}{x^{-}}^{j}-\langle{x^{+}}^{i}{x^{-}}^{j}\rangle\}_{i,j}$
@@ 7.4. Nonlinear drift
which can be proven by induction on $n$ [^20] The index “(1)” refers to $\mathbf{x}$ and is sometimes omitted, and the index “(0)” refers to the constant 1, so that $\langle\mathrm{d}\mathbf{x}^{(i)}(\tau)/\mathrm{d}\tau\mid\mathbf{x}(\tau)\rangle=\sum_{j=0}^{\infty}{\widehat{\mathbf{A}}}^{(i)}_{(j)}\mathbf{x}^{(j)}(\tau)$ and $\langle\mathbf{x}^{(i)}(\tau)\mid\mathbf{x}(0)\rangle=\sum_{j=0}^{\infty}(e^{\widehat{\mathbf{A}}\tau})^{(i)}_{(j)}\mathbf{x}^{(j)}(0)$.. Also:
@@ 7.5. Integrated variables
where the superscript “$y(i)$” refers to $\mathbf{x}^{(i)}\Delta y$ (“$y(0)$” being synonymous with “$y$”). Also, for $n\geq 1$, $i\geq 0$, $j\geq 0$:
@@ 8.3.1. Inference from data
$\langle x^{i}v^{j}v^{k}\rangle$, $\widetilde{L}(x^{i},v^{j},v^{k})$, $\langle v^{i}v^{j}v^{k}\rangle$, $L(v^{i}v^{j},v^{k})$ (satisfying Eq. (96)), $\langle x^{i}\mathrm{d}[v^{j},v^{k}]/\mathrm{d}t\rangle$, and $\langle v^{i}\mathrm{d}[v^{j},v^{k}]/\mathrm{d}t\rangle$. With the exception of the last two, these quantities can be straightforwardly estimated from time-lapse data without corrections due to fluctuations. The estimation of fluctuation covariance $\bm{\mathcal{B}}$ and measurement error covariance $\bm{\Lambda}$ is derived in the Supplementary Material in [10]. However, Eq. (S70) in [10] should read $(6/11)(-1,1,1,-3,1,1)^{\mathsf{T}}$ as in their code. We may write these estimators in terms of the measured values $\mathbf{y}:=\mathbf{x}+\bm{\eta}$, where $\bm{\eta}$ is measurement noise. Using the second differences:
(340) $\Delta^{2}\mathbf{y}_{(-)}$ $:=\mathbf{y}(t+2\Delta t)-2\mathbf{y}(t+\Delta t)+\mathbf{y}(t),$
(341) $\Delta^{2}\mathbf{y}_{(+)}$ $:=\mathbf{y}(t+3\Delta t)-2\mathbf{y}(t+2\Delta t)+\mathbf{y}(t+\Delta t),$
we may write the estimators as:
(342) ^$\widehat{\mathcal{B}}^{ij}$ $=\frac{3}{11(\Delta t)^{3}}\left({2\Delta^{2}y^{i}_{(-)}\Delta^{2}y^{j}_{(-)}+2\Delta^{2}y^{i}_{(+)}\Delta^{2}y^{j}_{(+)}+3\Delta^{2}y^{i}_{(-)}\Delta^{2}y^{j}_{(+)}+3\Delta^{2}y^{i}_{(+)}\Delta^{2}y^{j}_{(-)}}\right),$
(343) $\widehat{\Lambda}^{ij}$ $=\frac{1}{44}\left({\Delta^{2}y^{i}_{(-)}\Delta^{2}y^{j}_{(-)}+\Delta^{2}y^{i}_{(+)}\Delta^{2}y^{j}_{(+)}-4\Delta^{2}y^{i}_{(-)}\Delta^{2}y^{j}_{(+)}-4\Delta^{2}y^{i}_{(+)}\Delta^{2}y^{j}_{(-)}}\right).$
Unfortunately, the error in [10] is propagated to the subsequent calculations. Considering inhomogeneous diffusion, Eq. (S77) in [10] should vanish as a result of $\widehat{\mathcal{B}}$ being written in terms of second differences. The next-order correction is obtained by multiplying the force by the measurement error, in place of Eq. (S75). The resulting correction analogous to Eq. (S77) is of order $\mathcal{O}(\Lambda/(\Delta t)^{2})$, which can be ignored because the range of validity of the inference procedure is $\bm{\Lambda}\ll\langle\mathbf{v}\mathbf{v}^{\mathsf{T}}\rangle(\Delta t)^{2}$. Thus, the choice of position and velocity estimators in Eq. (S92) are of little consequence. The same is true for the position estimator in Eq. (S48). We may choose time-symmetric or -antisymmetric combinations to preserve time-symmetry or -antisymmetry. The error in Eq. (S48) should also be $\mathcal{O}(\Delta t,\Lambda/(\Delta t)^{2})$, where the $\mathcal{O}(\Lambda/(\Delta t)^{2})$ arises from second derivatives with respect to velocity of the basis functions.
@@ 8.3.2. Solution of third-order quantities
×$\quad{}\times\Big{[}{-(\mathbf{A}_{\mathbf{v}}^{-1})^{i^{\prime}}_{i^{\prime\prime}}\left({(\widetilde{a}_{xx})^{i^{\prime\prime}}_{lm}\langle\widetilde{x}^{j^{\prime}}\widetilde{x}^{l}\rangle\langle\widetilde{x}^{k^{\prime}}\widetilde{x}^{m}\rangle+\mu(a_{xv})^{i^{\prime\prime}}_{lm}\langle\widetilde{x}^{j^{\prime}}\widetilde{x}^{l}\rangle\langle\widetilde{x}^{k^{\prime}}\widetilde{v}^{m}\rangle+\mu(a_{xv})^{i^{\prime\prime}}_{lm}\langle\widetilde{x}^{j^{\prime}}\widetilde{v}^{m}\rangle\langle\widetilde{x}^{k^{\prime}}\widetilde{x}^{l}\rangle}\right)}\Big{.}$
@@
.$\qquad\Big{.}{{}+(\mathbf{A}_{\mathbf{v}}^{-1})^{j^{\prime}}_{j^{\prime\prime}}(\mathbf{A}_{\mathbf{v}}^{-1})^{k^{\prime}}_{k^{\prime\prime}}(b_{x})^{j^{\prime\prime}k^{\prime\prime}}_{l}\langle\widetilde{x}^{i^{\prime}}\widetilde{x}^{l}\rangle}\Big{]}+\mathcal{O}(\mu^{-1})(a_{vv},b_{v})+\mathcal{O}(\mu^{-2})(a_{xx},a_{xv},b_{x})$
@@
.$\qquad\Big{.}{{}-(\mathbf{A}_{\mathbf{v}}^{-1})^{i}_{i^{\prime}}(b_{x})^{i^{\prime}k^{\prime}}_{l}\langle\widetilde{x}^{j}\widetilde{x}^{l}\rangle-(\mathbf{A}_{\mathbf{v}}^{-1})^{j}_{j^{\prime}}(b_{x})^{j^{\prime}k^{\prime}}_{l}\langle\widetilde{x}^{i}\widetilde{x}^{l}\rangle}\Big{)}$
@@
×$\quad{}\times\Big{(}{(a_{vv})^{i^{\prime}}_{lm}\langle\widetilde{v}^{j^{\prime}}\widetilde{v}^{l}\rangle\langle\widetilde{v}^{k^{\prime}}\widetilde{v}^{m}\rangle+(a_{vv})^{j^{\prime}}_{lm}\langle\widetilde{v}^{i^{\prime}}\widetilde{v}^{l}\rangle\langle\widetilde{v}^{k^{\prime}}\widetilde{v}^{m}\rangle+(a_{vv})^{k^{\prime}}_{lm}\langle\widetilde{v}^{i^{\prime}}\widetilde{v}^{l}\rangle\langle\widetilde{v}^{j^{\prime}}\widetilde{v}^{m}\rangle}\Big{.}$
.$\qquad\Big{.}{{}+(b_{v})^{i^{\prime}j^{\prime}}_{l}\langle\widetilde{v}^{k^{\prime}}\widetilde{v}^{l}\rangle+(b_{v})^{i^{\prime}k^{\prime}}_{l}\langle\widetilde{v}^{j^{\prime}}\widetilde{v}^{l}\rangle+(b_{v})^{j^{\prime}k^{\prime}}_{l}\langle\widetilde{v}^{i^{\prime}}\widetilde{v}^{l}\rangle}\Big{)}$

### v17

@@ 4. Third-order properties
(100) $0$ $=-\int\mathrm{d}\mathbf{x}\,x^{i}x^{j}x^{k}\frac{\partial J^{l}(\mathbf{x})}{\partial x^{l}}$
@@ 5.2. Two-dimensional example
(124) $\lambda_{x}\langle xy^{4}\rangle=-4b^{xy}_{y}\langle y^{4}\rangle.$
@@ 5.3. Integrated variables
We note that if $D^{ky}=0$ for $k$ stationary, which is a condition for detailed balance, then $\langle x^{k}(\tau)(y(\tau)-y(0))\rangle=0$. If in addition $2b^{iy}_{k}C^{jk}+2b^{jy}_{k}C^{ik}=0$, then $\langle x^{i}(\tau)x^{j}(\tau)(y(\tau)-y(0))\rangle=0$.
@@ 6.2. One-dimensional case
(asterisk denoting the complex conjugate), which can be easily shown using Itô’s lemma (with suitable assumptions on $f,g$). When $a=b=0$, the Koopman eigenfunctions are the probabilist’s Hermite polynomials ${He}_{n}(x)$ with eigenvalues n-n. In the present case, the Koopman eigenfunctions $f_{n}(x)$ and the corresponding eigenvalues $\lambda_{n}$ are:
@@ 6.3. Multidimensional case
$\qquad{}\times\left({a^{i^{\prime}}_{lm}C^{j^{\prime}l}C^{k^{\prime}m}+a^{j^{\prime}}_{lm}C^{i^{\prime}l}C^{k^{\prime}m}+a^{k^{\prime}}_{lm}C^{i^{\prime}l}C^{j^{\prime}m}+b^{i^{\prime}j^{\prime}}_{l}C^{k^{\prime}l}+b^{i^{\prime}k^{\prime}}_{l}C^{j^{\prime}l}+b^{j^{\prime}k^{\prime}}_{l}C^{i^{\prime}l}}\right)$
@@ 6.4. Generalized Koopman eigenfunctions
$\gamma^{ij}_{klm}$ $:=\left\{{\left[{(\mathbf{A}\otimes\mathds{1}+\mathds{1}\otimes\mathbf{A})\otimes\mathds{1}\otimes\mathds{1}\otimes\mathds{1}-\mathds{1}\otimes\mathds{1}\otimes(\mathbf{A}\otimes\mathds{1}\otimes\mathds{1}+\mathds{1}\otimes\mathbf{A}\otimes\mathds{1}+\mathds{1}\otimes\mathds{1}\otimes\mathbf{A})}\right]^{-1}}\right\}^{ijk^{\prime}l^{\prime}m^{\prime}}_{i^{\prime}j^{\prime}klm}$
(173) $\quad{}\times\left({a^{i^{\prime}}_{k^{\prime}l^{\prime}}\delta^{j^{\prime}}_{m^{\prime}}+a^{j^{\prime}}_{k^{\prime}l^{\prime}}\delta^{i^{\prime}}_{m^{\prime}}}\right).$
@@
$\qquad{}\times\left({x^{k^{\prime}}(0)x^{l^{\prime}}(0)x^{m^{\prime}}(0)-C^{k^{\prime}l^{\prime}}x^{m^{\prime}}(0)-C^{k^{\prime}m^{\prime}}x^{l^{\prime}}(0)-C^{l^{\prime}m^{\prime}}x^{k^{\prime}}(0)}\right)$
@@ 6.7. Finite time interval inhomogeneous diffusion
However, for finite $\tau$, the above quantity can be non-zero for homogeneous diffusion because of non-vanishing third moments. This can be illustrated by the example of Eq. (93) in one dimension with A=D=1-A=D=1 and $b=0$. From Eqs. (156)–(157), we have:
@@ 6.8. Quantitative significance of third-order effects
$\qquad\qquad\qquad{}\times\left({D^{yz}-A^{y}_{p}(\mathbf{A}^{-1})^{p}_{q}D^{qz}-A^{z}_{r}(\mathbf{A}^{-1})^{r}_{s}D^{sy}+A^{y}_{p}A^{z}_{r}(\mathbf{A}^{-1})^{p}_{q}(\mathbf{A}^{-1})^{r}_{s}D^{qs}}\right).$
@@
$\qquad\Big.{{}+C^{ji^{\prime}}(D^{ij^{\prime}}D^{kk^{\prime}}-D^{ik^{\prime}}D^{kj^{\prime}})+C^{jj^{\prime}}(D^{ii^{\prime}}D^{kk^{\prime}}-D^{ik^{\prime}}D^{ki^{\prime}})}\Big].$
@@ 6.10. Comparison of experimental data and theoretical predictions
We consider the sum and difference of these quantities at $\tau=0$. The former is equal to $(\Delta t)L(x^{i}x^{j},y)$, while the latter is equal to $2(\Delta t)(C^{ik}b^{jy}_{k}+C^{jk}b^{iy}_{k})$. We can calculate the “ensemble variances” of these quantities using the aforementioned prescriptions. We may consider using them for comparison at arbitrary $\tau$. However, we may be interested in comparing the above quantities directly. They may be obtained by taking one-half the sum and difference of the sum and difference. We may treat the sum and difference of the above quantities as uncorrelated in the “ensemble covariance” as they acquire opposite signs under time reversal. Thus we conclude that the above quantities should be compared according to one-fourth the sum of the “ensemble variances” of $(\Delta t)L(x^{i}x^{j},y)$ and $2(\Delta t)(C^{ik}b^{jy}_{k}+C^{jk}b^{iy}_{k})$. As in the linear case, and as is analogously true for all quantities, if the theoretical values at $\tau=0$ exceed the prescriptions of the “ensemble variances” (for either second- or third-order quantities), we may need to increase the tolerance for comparison accordingly.
@@ 7.2. Detailed balance
Here we discuss the conditions for detailed balance in the Fokker–Planck equation Eq. (97) in the presence of odd variables. Each variable $x^{i}$ has a parity $\epsilon^{i}=\pm 1$, where under time-reversal $x^{i}\to\epsilon^{i}x^{i}$ (ii not summed over), and the product $\bm{\epsilon}\mathbf{x}$ is defined as $(\bm{\epsilon}\mathbf{x})^{i}=\epsilon^{i}x^{i}$ (ii not summed over), following [3]. In terms of the stationary distribution $p(\mathbf{x})$, the conditions for detailed balance are [3]:
@@ 7.3. Inhomogeneous diffusion
(248) $\mathbf{x}^{(2)+}$ $:=\{{x^{+}}^{i}{x^{+}}^{j}-\langle{x^{+}}^{i}{x^{+}}^{j}\rangle,{x^{-}}^{i}{x^{-}}^{j}-\langle{x^{-}}^{i}{x^{-}}^{j}\rangle\}_{i,j}$
(249) $\mathbf{x}^{(2)-}$ $:=\{{x^{+}}^{i}{x^{-}}^{j}-\langle{x^{+}}^{i}{x^{-}}^{j}\rangle\}_{i,j}$
@@ 7.4. Nonlinear drift
which can be proven by induction on nn [^20] The index “(1)” refers to $\mathbf{x}$ and is sometimes omitted, and the index “(0)” refers to the constant 1, so that $\langle\mathrm{d}\mathbf{x}^{(i)}(\tau)/\mathrm{d}\tau\mid\mathbf{x}(\tau)\rangle=\sum_{j=0}^{\infty}{\widehat{\mathbf{A}}}^{(i)}_{(j)}\mathbf{x}^{(j)}(\tau)$ and $\langle\mathbf{x}^{(i)}(\tau)\mid\mathbf{x}(0)\rangle=\sum_{j=0}^{\infty}(e^{\widehat{\mathbf{A}}\tau})^{(i)}_{(j)}\mathbf{x}^{(j)}(0)$.. Also:
@@ 7.5. Integrated variables
where the superscript “y(i)y(i)” refers to $\mathbf{x}^{(i)}\Delta y$ (“y(0)y(0)” being synonymous with “yy”). Also, for $n\geq 1$, $i\geq 0$, $j\geq 0$:
@@ 8.3.1. Inference from data
[…]
@@ 8.3.2. Solution of third-order quantities
$\quad{}\times\Big[{-(\mathbf{A}_{\mathbf{v}}^{-1})^{i^{\prime}}_{i^{\prime\prime}}\left({(\widetilde{a}_{xx})^{i^{\prime\prime}}_{lm}\langle\widetilde{x}^{j^{\prime}}\widetilde{x}^{l}\rangle\langle\widetilde{x}^{k^{\prime}}\widetilde{x}^{m}\rangle+\mu(a_{xv})^{i^{\prime\prime}}_{lm}\langle\widetilde{x}^{j^{\prime}}\widetilde{x}^{l}\rangle\langle\widetilde{x}^{k^{\prime}}\widetilde{v}^{m}\rangle+\mu(a_{xv})^{i^{\prime\prime}}_{lm}\langle\widetilde{x}^{j^{\prime}}\widetilde{v}^{m}\rangle\langle\widetilde{x}^{k^{\prime}}\widetilde{x}^{l}\rangle}\right)}\Big$.
@@
$\qquad\Big.{{}+(\mathbf{A}_{\mathbf{v}}^{-1})^{j^{\prime}}_{j^{\prime\prime}}(\mathbf{A}_{\mathbf{v}}^{-1})^{k^{\prime}}_{k^{\prime\prime}}(b_{x})^{j^{\prime\prime}k^{\prime\prime}}_{l}\langle\widetilde{x}^{i^{\prime}}\widetilde{x}^{l}\rangle}\Big]+\mathcal{O}(\mu^{-1})(a_{vv},b_{v})+\mathcal{O}(\mu^{-2})(a_{xx},a_{xv},b_{x})$
@@
$\qquad\Big.{{}-(\mathbf{A}_{\mathbf{v}}^{-1})^{i}_{i^{\prime}}(b_{x})^{i^{\prime}k^{\prime}}_{l}\langle\widetilde{x}^{j}\widetilde{x}^{l}\rangle-(\mathbf{A}_{\mathbf{v}}^{-1})^{j}_{j^{\prime}}(b_{x})^{j^{\prime}k^{\prime}}_{l}\langle\widetilde{x}^{i}\widetilde{x}^{l}\rangle}\Big)$
@@
$\quad{}\times\Big({(a_{vv})^{i^{\prime}}_{lm}\langle\widetilde{v}^{j^{\prime}}\widetilde{v}^{l}\rangle\langle\widetilde{v}^{k^{\prime}}\widetilde{v}^{m}\rangle+(a_{vv})^{j^{\prime}}_{lm}\langle\widetilde{v}^{i^{\prime}}\widetilde{v}^{l}\rangle\langle\widetilde{v}^{k^{\prime}}\widetilde{v}^{m}\rangle+(a_{vv})^{k^{\prime}}_{lm}\langle\widetilde{v}^{i^{\prime}}\widetilde{v}^{l}\rangle\langle\widetilde{v}^{j^{\prime}}\widetilde{v}^{m}\rangle}\Big$.
$\qquad\Big.{{}+(b_{v})^{i^{\prime}j^{\prime}}_{l}\langle\widetilde{v}^{k^{\prime}}\widetilde{v}^{l}\rangle+(b_{v})^{i^{\prime}k^{\prime}}_{l}\langle\widetilde{v}^{j^{\prime}}\widetilde{v}^{l}\rangle+(b_{v})^{j^{\prime}k^{\prime}}_{l}\langle\widetilde{v}^{i^{\prime}}\widetilde{v}^{l}\rangle}\Big)$